4/7 of a pole in a mud . when 1/4 of this part is pulled out ,5/2 meters of the pole is still in the mud . what is the full length of the pole?
Answer by mananth(16946)

let length of pole be x
4x/7 is in mud
1/4 of 4x/7 is pulled out
4x/28 is out
..
5/2 is still in the mud
..
x-4x/28= 5/2
28x-4x / 28 = 5/2
2(24x)=28*5
48x=140
x=2.91 m the length of the pole